You plan to deposit $100 each month into an IRA earning 0.40% interest monthly. How much will you have in your account in 15 years?
Answer #1

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

n = 15 * 12 = 180 months

(1+0.004) 180 – 1 FV = 100* [ 0.004

FV = 100 * 262.8712

FV = $26,287.12